(a ). You rec only 3.0%. what is the rate of inflation? eived a raise in salary of 8%. However, your purchasing power went up (b). (Unrelated to (a). What is the CPI?
(c). Your company wishes to fund a new project by raising funds through three sources: Common Stock. $5 M, Retained Earning. $6M. and a bank loan. 10 M. The income tax rate is .35. The interest rate on the loan is 6% compounded quarterly. Earning per share are 0.75 dollars per year with the Price per Share taken to be 11.00 dollars. What is the Weighted Average Cost of Capital?
Answer #1

increase in nominal income = increase in real income + inflation
8% = 3% + inflation
inflation = 5%

CPI measures the consumer price index. It is an index that tracks the weighted average prices of a representative basket of goods and services. CPI = rate of inflation = 5%

cost of debt = (1+6%/4)^4-1 = 6.14%
weight of debt = 10/21
cost of equity = 0.75/11 = 6.82%
weight of equity = 11/21

WACC = weight of debt*cost of debt + weight of equity*cost of equity
WACC = 10/21*6.14*(1-35%) + 11/21*6.82% = 5.47%